OpenStudy (anonymous):

The product of the page number on two facing pages of a book is 306. Find the page numbers?

OpenStudy (amistre64):

306+1 ------ maybe? 2

OpenStudy (amistre64):

306/2 = 153 .... so thats kinda tricky

OpenStudy (amistre64):

i dont think it can exist; 1,5;4,2;3,3 are your option to =6

OpenStudy (amistre64):

ack..gotta learn to read

OpenStudy (amistre64):

mn = 306 lol

OpenStudy (anonymous):

so that is the answer sir

OpenStudy (amistre64):

no, its my other voices correcting me :)

OpenStudy (anonymous):

o ok well you let me which I need to put I have this question and one more that are posted

OpenStudy (amistre64):

10*10 = 100 25*26 = nothing good

OpenStudy (amistre64):

8*7 = 56; so that possible endings

OpenStudy (amistre64):

18*17?

OpenStudy (amistre64):

yep; 18*17 = 306

OpenStudy (amistre64):

20 was to high; 15 to low, so that was my only logical option to test :)

OpenStudy (anonymous):

17,18 is correct

OpenStudy (amistre64):

yay!! it was either that or start using imaginary numbers :P

OpenStudy (gw2011):

I also got 17 and 18 by using the quadratic formula.

OpenStudy (anonymous):

i got it by taking the square root of 306 and then taking the nearest integer on either side